OpenStudy (anonymous):

what is the solution set for this inequality? x/4+x/5

OpenStudy (anonymous):

Tell me if I'm right? x is less than or equal to 20? thats what i got

hero (hero):

Okay, so here's what you do:

hero (hero):

\(\frac{x}{4} + \frac{x}{5} \le \frac{9}{1}\)

OpenStudy (anonymous):

ok

hero (hero):

Multiply 9/1 by 5/5 to get: \(\frac{x}{4} + \frac{x}{5} \le \frac{9}{1} \dot\ \frac{5}{5}\)

hero (hero):

Simplify that, then subtract x/5 from both sides

OpenStudy (anonymous):

Is it x/4=5

OpenStudy (anonymous):

<**

hero (hero):

No, you have to multiply the right side

hero (hero):

You know how to multiply fractions right?

OpenStudy (anonymous):

i multiplied it and got 5

hero (hero):

You multiplied \(\frac{9}{1} \dot\ \frac{5}{5}\) and got 5?

OpenStudy (anonymous):

no 9

hero (hero):

The point is to multiply and get 45/5

hero (hero):

you need two fractions with the same denominator in order for this to work

hero (hero):

When you subtract x/5 from both sides you should get: \(\frac{x}{4} = \frac{45}{5} - \frac{x}{5}\)

OpenStudy (anonymous):

is my answer right?

OpenStudy (anonymous):

that was my answer. x is less than or equal to 20

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

@Hero

hero (hero):

yes

hero (hero):

I was trying to show you how to do it without needed to find an lcd. All you have to do is get two fractions with the same denominator then combine them together

OpenStudy (anonymous):

ok!

hero (hero):

Trust me, there will be more complicated fractional equations and inequalities you will run into, and knowing an alternative method will help